OpenStudy (anonymous):

HELPP!!! Identify the horizontal asymptote of f(x) = quantity 6 x minus 7 over quantity 11 x plus 8. y = 0 y = 6 over 11 y = negative 7 over 8 No horizontal asymptote

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ f(x)= \frac{6x-7}{11x+8} }\)

OpenStudy (solomonzelman):

Lets make an observation. at first, you know that \(\LARGE\color{black}{ 11x+8\neq0 }\) and this value that x can't equal is the vertical asymptote.

OpenStudy (solomonzelman):

Can you find the vertical asymptote?

OpenStudy (anonymous):

idk how which is why i mainly need help with this

OpenStudy (solomonzelman):

the denominator can't equal zero, you know this, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because we can never divide by 0

OpenStudy (solomonzelman):

right

OpenStudy (solomonzelman):

So, this x-value, where the denominator is zero, is where the function doesn't have any value. It is undefined there.

OpenStudy (solomonzelman):

Okay?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

lets find this value. 11x+8=0 (where does the function NOT exist) subtracting 8 from both sides 11x=-8 dividing by 11 on both sides, x=-8/11 this is the vertical asymptote.

OpenStudy (solomonzelman):

good so far?

OpenStudy (anonymous):

yes :)

OpenStudy (solomonzelman):

I'll tell you hat does it mean to have a vertical asymptote.

OpenStudy (solomonzelman):

is

OpenStudy (anonymous):

okay :)

OpenStudy (solomonzelman):

If you try to hit an x-value that is close to this x=-8/11 will approximate it to -0.73. so, our asymptote is x=-0.73 if you try x=-0.72 you will get a value that is a little bit greater than zero. So the closer you are getting to -0.73 (the -8/11) the smaller the denominator will be right?

OpenStudy (solomonzelman):

This means that as x approaches -8/11, the y gets infinitely large. (keep in mind, it doesn't rich exactly -8/11, because there is no value there)

OpenStudy (solomonzelman):

I'll give you time to process this.

OpenStudy (solomonzelman):

ask any why questions if you don't get something in the middle of the text....

OpenStudy (anonymous):

right i know that the closer we're getting to -8/11 the smaller the denominator will be

OpenStudy (solomonzelman):

yes, so a very small positive decimal (like 0.00001) will make the f(x) 'S value infinitely large. this is when we choose an x value that is a tiny bit more than -8/11. (by more I keep in mind that -4<-3 or -2<-1... you know what I mean? a larger negative number like -5, is SMALLER than a small negative number like -4)

OpenStudy (solomonzelman):

So from the right, as x gets closer to -8/11, the y goes infinitely up and up.

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

BUT, the top is negative. because the top is "6x-7" so when we plug in something close to -8/11 (or any negative value) for x, the numerator is negative. So... For numbers that are a big greater than -8/11, while we get a small positive decimal on the bottom, on the top though we get a fairly big negative number. And big negative number ----------------- = NEGATIVE INFINITY (ok ?) small positive decimal

OpenStudy (anonymous):

right

OpenStudy (solomonzelman):

Okay. We have observed that. (This is actually a calculus-wise understanding of an asymptote)

OpenStudy (anonymous):

LOL wooow

OpenStudy (solomonzelman):

Now, lets go from the other side of the asymptote. We have plugged in numbers that are on the right of the asymptote. And this gave us a picture like,